OpenStudy (anonymous):

Write an equation to describe the given situation: Paula's purse contains twice as many dimes as quarters, and three fewer nickels than dimes. The total of these coins is $3.15.

OpenStudy (anonymous):

Set each type of coin as a variable: d=# of dimes, q=# of quarters, n=# of nickels. Now, given relationships between the numbers of the coins in equation form: d=2q or q=d/2, n=d-3. Now create an expression that only uses one variable, d to represent the total number of coins: d+d/2+d-3. Now, to represent the value of each type of coin, multiply each coin by its respective value: d(0.10)+d/2(0.25)+(d-3)(0.05). Now set the expression equal to the total value of the coins, and you have your equation